An object is thrown vertically upward with an initial velocity of 75ft/sec from a building 40ft above the ground. (a = 32ft/sec²)


1. Determine the equation of the distance s from the ground at a given time t in seconds.


2. Calculate the distance of the object from the ground after 2 seconds.



1
Expert's answer
2021-06-18T09:21:43-0400

1. Let y-axis is directed upward from the ground. Then "y(t)" will be the distance from the ground at a given time "t" in seconds.

"a=-32\ ft/s^2, v_0=75\ ft/s, y_0=40\ ft"

"v(t)=\int a dt=v_0+at"

"y(t)=\int vdt=y_0+v_0t+\dfrac{at^2}{2}"

The equation of the motion along y-axis is


"y(t)=y_0+v_0t+\dfrac{at^2}{2}, t\geq0"

Substitute


"y(t)=40+75t-\dfrac{32t^2}{2}, ft"

"y(t)=40+75t-16t^2, ft"

2.


"y(2)=40+75(2)-16(2)^2=126(ft)"

The distance of the object from the ground after 2 seconds is 126 ft.